OpenStudy (anonymous):

Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. A (n) = -3 + (n – 1)(-2.2) A) -3, -11.8, -25 B) 0, -6.6, -19.8 C) -3, -9.6, -22.8 D) -2.2, -11.8, -19.8

OpenStudy (anonymous):

@studygeek15

OpenStudy (anonymous):

Just put in 1, 4 and 10 one by one instead of n

OpenStudy (anonymous):

A (1) = -3 + (1 - 1)(-2.2) @ArsalanAijaz

OpenStudy (anonymous):

Do what @ArsalanAijaz said. For the first one A (1) = -3 + (1 – 1)(-2.2) and solve.

OpenStudy (anonymous):

Aye sir :P

OpenStudy (anonymous):

No for the first one. A (1) = -3 + (1 – 1)(-2.2) = -3 + (0)(-2.2) = -3 + 0 = -3 Remember that anything times 0 is equal to 0

OpenStudy (anonymous):

Do the same thing for 4. Let me know what number you get.

OpenStudy (anonymous):

A (n) = -3 + (n – 1)(-2.2) A (4) = -3 + (4-1)(-2.2) = -3 + (3)(-2.2) = -3 + 6.6 = 9.6

OpenStudy (anonymous):

Did I do that correct? @studygeek15

OpenStudy (anonymous):

You have the right idea. 3 * -2.2 is -6.6 and add -3. Your number is correct, but you need a negative sign.

OpenStudy (anonymous):

Oh okay, thank you

OpenStudy (anonymous):

No problem.